Two positive numbers have a difference of 1. The larger number is 4 less than 2 times the smaller number. What are the two numbers. I have no clue how to set this up and I need to help my daughter, who has a learning disability, understand her math. I would greatly appreciate it if someone could please set the problem up for me and possibly explain why the steps that are taken are done so I can explain it to her.
Answer by ReadingBoosters(3246)

Start by defining your variables
x = larger integer
y = smaller integer
...
Difference is 1, so subtract them
x-y = 1
...
x is 4 less (minus) than 2 times (multiply by two) the smaller number, y
x=2y-4
...
Substitute x into the first equation
(2y-4) - y =1
...
Combine like terms
2y - y -4 = 1
y - 4 = 1
Add 4 to both sides
y = 5
...
The small integer is 5 and the larger integer is
x=2y-4
x=2(5)-4 = 10-4 = 6
The larger integer is 6
...
Proof
6 - 5 = 1
...............................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.